OpenStudy (anonymous):

I WILL FAN AND MEDAL Simplify completely: 14x^5y^4+21x^3y^2/7x^3y The / means over (like a fraction

OpenStudy (anonymous):

\[\frac{ 14x^5y^4+21x^3y^2 }{ 7x^3y}\]

OpenStudy (anonymous):

2x8y5 + 3x6y3 2x2y4 + 3xy 2x2y3 + 3y 2x2y3 + 3y2

OpenStudy (anonymous):

thw simplified form is : \[2x^2y^3+3y\]

OpenStudy (fifciol):

tip: \[\frac{ a+b }{c }=\frac{ a }{ c }+\frac{ b }{ c }\]

OpenStudy (anonymous):

you just need to take a factor of 7x^3y from the numerator and then it can be canceled with denominator .